OpenStudy (anonymous):

Assume a spinner has four unequal sectors red (0.33), blue (0.27), green (0.18),and white (0.22). If the spinner is spun twice, find the probability that the spinner will land on different colors both times.

OpenStudy (kropot72):

On the first spin say spinner lands on red. On the second spin P(not red) = 1 - P(red) = 1 - 0.3 The overall probability of different colors both times = 0.33(1 - 0.33) Do the same for blue, green and white as the first color, then sum the four results to find the overall probability that the spinner will land on different colors both times.